OpenStudy (anonymous):

derivative of (1-t)(1+t^2)^-1

OpenStudy (phi):

do you know the product rule for taking the derivative of two expressions that are multiplied?

OpenStudy (anonymous):

do you know this method ? u/v=u'v-v'u/v^2 ?

OpenStudy (anonymous):

phi i think the division rule is simpler for him since the 2nd product is ^-1 in which simply you can put it in the deniminator

OpenStudy (anonymous):

i do know the product rule, i was just getting lost in the middle of the madness

OpenStudy (phi):

let u = (1-t) and v= (1+t^2)^-1 du/dt = -1 dv/dt uses the exponent rule and the chain rule: \[ \frac{d}{dt}\left( 1+t^2\right)^{-1} = -1 \cdot \left( 1+t^2\right)^{-1-1} \frac{d}{dt}\left( 1+t^2\right)\] can you finish finding dv/dt ? then put the pieces together d(u v) = u dv + v du